COMPOUND INTERESTI
Find the future value (compound amount) and the compound interest, as indicated, for
each of the following investments. Round answers to the nearest cent. Use a calculator
or Table 16-1 to find FVF.
Present
Term
Value
Rate
4 years
.S6,000
6% compounded monthly
2. $960
6% compounded semiannually
years
8 years
3.
S20,000
8% compounded quarterly
1
Expert's answer
2020-12-11T06:02:15-0500

Future value formula:

FV = PV* (1+r)n; FV - future value, PV - present value, r - rate of return, n - number of periods

compound interest = FV-PV

1) PV = 6000

n = 12*4 = 48

r = 0.06/12 = 0.005

FV = 6000*(1+0.005)48 = 7622.93

compound interest = 7622.93-6000 = 1622.93

2) PV = 960

n = 11*2 = 22

r = 0.06/2 = 0.03

FV = 960*(1+0.03)22 = 1839.46

compound interest = 1839.46-960 = 879.46

3) PV = 20000

n = 8*4 = 32

r = 0.08/4 = 0.02

FV = 20000*(1+0.02)32 = 37690.80

compound interest = 37690.80-20000 = 17690.80